OpenStudy (anonymous):

help!!! simplify do not use negative exponents in the answer 12a 5th power / 2a 8th power

OpenStudy (anonymous):

\[12a ^{5}\2a ^{8}\] real problem

OpenStudy (anonymous):

Just to be clear, is this what you mean? \[\frac{12a^{5}}{2a^{8}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, are you able to solve it *with* negative exponents?

OpenStudy (anonymous):

I did not know how to make the line on here

OpenStudy (anonymous):

\frac{top}{bottom} Do that in the Equation editor, or within \[ and \ ]

OpenStudy (anonymous):

nope it said do not use neg exponents and the 5 is -5 and the 8 is -8

OpenStudy (anonymous):

I know, but once we get to the final answer *with* negative exponents, we can get rid of them by re-writing the answer.

OpenStudy (anonymous):

Wait, are you saying it's to the power of -5 and -8?

OpenStudy (anonymous):

Which is correct—top or bottom? \[\frac{12a^{5}}{2a^{8}}\] \[\frac{12a^{-5}}{2a^{-8}}\]

OpenStudy (anonymous):

bottom

OpenStudy (anonymous):

Okay, great. And there are no brackets, right? It's not \((12a)^{-5}\)?

OpenStudy (anonymous):

So, to start, are you able to simplify the equation?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I could not find an exmaple in the book like mines so one of the problems in the exercise is like my real problem i need help to figure it out so i can solve mines

OpenStudy (anonymous):

Do you know how to simplify exponents?

OpenStudy (anonymous):

When dividing like bases (in this case, \(a\)), you subtract the bottom exponent from the top. So, \(a^{(-5)-(-8)}\).

OpenStudy (anonymous):

As for your numeric constants, you just divide them (\(12 \div 2\))

OpenStudy (anonymous):

ok so that will be -13

OpenStudy (anonymous):

5-8

OpenStudy (anonymous):

No, it's (-5) - (-8).

OpenStudy (anonymous):

12/2 is 6 -5-(-8)=3

OpenStudy (anonymous):

A double-negative is a positive (in math).

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

So what's your final answer?

OpenStudy (anonymous):

6a\[6a ^3\]

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

do I do the samething if my problem has 12 x top and 8y bottom

OpenStudy (anonymous):

Do you mean like: \(\Large{\frac{12x}{8y}}\) ? No, those are different variables, so you cannot divide one by the other.

OpenStudy (anonymous):

You can simplify the numbers \(\Large{\frac{12}{8} = \frac{3}{2}}\), but that's all.

OpenStudy (anonymous):

yes with -6 at the top and -10 at the bottom

OpenStudy (anonymous):

Ahh, so: \[\frac{12x^{-6}}{8y^{-10}}\] Nope, all you can simplify in this case is 12/8.

OpenStudy (anonymous):

3x,2y

OpenStudy (anonymous):

Yes, with the same exponents.

OpenStudy (anonymous):

so there isn't any steps to break it down because my teacher keep saying show your work

OpenStudy (anonymous):

If they're different variables, you can't simplify them further.

OpenStudy (anonymous):

The only "work" you could show is dividing both the top and bottom by 4, to simplify your numeric fraction.

OpenStudy (anonymous):

\[\Large\frac{12x^{-6}}{8y^{-10}} \small\frac{\div 4}{\div 4}\]

OpenStudy (anonymous):

ok